### v3

@@ 4.1 Data collection
In our research, we followed a multi-criteria approach that balanced geographic representation, data availability, and cultural diversity while working within practical constraints. Our selection methodology employed a three-tier filtering process: (1) continental representation to ensure global coverage, (2) availability in established datasets (MIT PP2 [6] and NUS Global Streetscapes [52]) to enable comparative analyses, and (3) consideration of cultural, urban form, and socioeconomic diversity within these constraints. This systematic approach reduced our candidate cities considerably (Global Streetscapes includes 688 cities around the world, and PP2 only covers 56) but ensured scientific rigor and comparability with existing literature. To ensure geographic diversity (1), we included the top five largest continents by population: South America, Europe, Africa, Asia, and North America. In terms of data availability (2), in South America, only three and two cities from Brazil and Chile are present in PP2, respectively. We chose Santiago as it was the only capital among the two countries listed in both PP2 and Global Streetscapes datasets, and a better cultural representation of a Hispanic city in South America. On one hand, while the number of candidate cities in Europe, Asia, and North America is higher, we opted for cities that are commonly used in perception studies due to their unique urban form and cultural characteristics [34, 43] (3): Amsterdam, Singapore, and San Francisco. On the other hand, only a handful of cities from South Africa and Botswana in Africa are present in PP2. While we aimed to prioritize the overlap between the cities in PP2 and the Global Streetscapes datasets, we did not consider Cape Town and Johannesburg, in South Africa, as representative candidates for the continent. Unfortunately, the other candidate city, Gaborone in Botswana, presented logistical difficulties when recruiting participants. Therefore, among the top five countries based on GDP in the continent (South Africa, Egypt, Algeria, Nigeria, and Morocco) [53], we opted for Abuja in Nigeria purely based on its central location in the continent and its presence in the Global Streetscapes. This city selection, despite its limitations, provides sufficient diversity in urban morphology, cultural contexts, and socioeconomic conditions (Figure 6) to test our core hypothesis about demographic influences on urban perception. We followed the image download process detailed in [52] and downloaded all available street view images within a 2.4 km2 square in the city center. For the city of Abuja, Nigeria, we manually selected a 2.4 km2 region with more urban development since the images in the Global Streetscapes dataset contained mostly highways scenes.
@@
Based on the contextual metadata available in the dataset, we selected only images with good quality, no reflection, non-panoramic, captured in clear weather, and taken during the day with front or back view direction. For image validity screening, we used visual complexity, as suggested in the Global Streetscapes dataset [52]. A low visual complexity value means the image is composed of predominantly one element, e.g., a road or wall. Fig. 7 shows the distribution of visual complexity values for all filtered images grouped by city. Images from all cities but Abuja have a clearer overlap in their distribution, with a visual complexity mean of 1.7; thus, we chose a conservative threshold of 1.5 and retained all images with at least this visual complexity value. However, images from Abuja have lower visual complexity values, with a mean of 1.3. While this overall lower visual complexity may reflect the intrinsic urban landscape of the selected region in the city, we aimed to have images with similar characteristics across all cities. Therefore, we retained images with at least a visual complexity of 1.3 as a trade-off for image diversity; there are twice the number of images with a visual complexity $\geq 1.3$ than with a visual complexity $\geq 1.5$.
@@ 4.2.1 Demographic survey
Participants were recruited from the general population, following their national representation by age (adults over 21 years) and gender for each country, with soft quotas for income level. The latter means that for the greater Amsterdam area, Singapore, and San Francisco and Santa Clara, participants were recruited such that their annual income ranges are close in quantity. For Chile and Nigeria, participants were recruited on a natural fallout basis. The available answers for the annual income ranges were adjusted to each country’s currency and situation based on the vendor’s suggestions, and the entire survey was translated into the official language in each country where it was deployed. Participants who did not reside in either of the five countries were excluded. A complete list of questions and their possible answers is listed in  subsection B.1.
@@ 4.2.2 Human perception rating
In the online survey, each participant rated five random unique pairs of images per indicator, totaling 50 pairwise comparisons across all 10 indicators. We chose the pairwise comparison scoring method because it has been shown to produce more stable results in urban visual perception studies [34]. Fig. 8 shows a screenshot of the online survey for one pairwise comparison under the lively indicator. Participants clicked on the image they preferred as an answer to the indicator question, e.g., the question “Which place looks livelier” in the screenshot, or chose “Both are the same to me” for an equally rated answer. Participants were not given a predetermined definition of the perceptual indicators. Instead, we aimed to captured their broad subjective interpretation of the ten different indicators. The complete perception survey questions are listed in subsection B.2.
@@ 4.3 Dataset
Our final dataset consists of images showcasing urban scenes from the cities of Santiago, Amsterdam, Abuja, Singapore, and San Francisco. A total of 400 images, 80 per city, were collected, and 200 respondents from each city (or neighboring areas) were recruited as raters, adding to 1,000 respondents in total. After deploying the online survey for two months, each participant answered a total of 61 questions, consisting of 11 demographic questions and 50 pairwise comparisons. We name this dataset Street Perception Evaluation Considering Socioeconomics (SPECS) and release it openly. Table 1 shows a summary of our collected dataset in the reporting format proposed by [34] and Table 3 in Appendix C shows the demographic background of the recruited participants from each country.
@@
Raters Demographics, sample size, number of ratings per image, recruitment process We recruited 1,000 respondents (200 from each country) who reside in the city, or neighboring cities, where the images were extracted – Santiago, greater Amsterdam area, Nigeria, Singapore, and San Francisco and Santa Clara – and are above 21 years old. An equal distribution of respondents, based on annual income range, was prioritised for all but Chile and Nigeria. The gender ratio is 1:1 within countries, and the age group ratio is roughly 1:1:1:2 for 21-29, 30-39, 40-49, and above 50. More than 78% of participants from each country have lived there for more than five years, and less than 5% have lived there for less than a year. Background information from all participants is shown in Table 3 in Appendix C. Each participant completed 50 ratings, with each image receiving an average of 25 ratings (min. 10, max. 46) per indicator.
Scoring Scoring methodology, rationale for method selection, evaluative indicators We selected pairwise comparison as the ranking scoring method due to its proven stability over and fairly quick assessment. On the online survey, participants were asked to choose their preference between two images given an indicator. To avoid a forced choice scenario that could lead to bias [68], and to be consistent with previous and future studies, the option for an equally rated preference (i.e., “Both are the same to me”) was shown too. We asked participants to rate a random pair of images five times for each of the 10 indicators, making up the 50 ratings per participant. Fig. 8 shows a screenshot of the online survey.
@@ 4.4 Data analysis
We calculated the perception scores using the Strength of Schedule (SOS) method, first used in urban perception in the Place Pulse 1.0 dataset [9]. This method’s score is referred to as the Q score, and it reflects the magnitude of the perception, bounded in $[0,10]$. Another commonly used method is the TrueSkill score [69], but unlike Q scores, it requires a higher minimum number of pairwise comparisons; i.e., 22 and 29, respectively; to reach stable scores [34]. Details about the Q score calculation are found in Appendix D. Relative scores such as Q scores and TrueSkill scores inherently depend on the ratings considered for their calculations. Thus, rather than being an absolute measure of an image’s perceptual dimensions, these scores are only comparative; they are only valid within the rating pool in which they were calculated. As shown in the bottom panel of Figure 1, step 3, we also included the predicted perception scores for all images using an existing perception model. This perception model, a Vision Transformer (ViT) developed by [70] and used in the Global Streetscapes dataset [52], serves as a proxy of a global or one-size-fits-all approach as it was fine-tuned with the PP2 dataset [6]. Specifically, the model was trained on the top and bottom 5% rated images for each perception dimension, which were selected as the representative positive and negative samples, respectively, adapting the method by [5]. The model was then trained, with a 80%-to-20% train-test split, to predict the probability of an image being assigned to the positive label, which was subsequently multiplied by 10 to obtain the perception score. This model, which we refer to as ViT-PP2, outputs perception scores for the six traditional indicators of the PP2 dataset and reported a higher accuracy in all but the boring indicator compared to the model proposed with the PP2 dataset.
@@
Lower 20,400,000 CLP & below 29,999 euros & below 2,000,000 NGN & below SG\$0 - SG\$44,999 US\$0 - US\$49,999
Middle 20,400,001 CLP - 60,400,000 CLP 30,000 euros - 81,999 euros 2,000,001 NGN- 4,000,000 NGN SG\$45,000 - SG\$99,999 US\$50,000 - US\$99,999
Upper 60,400,001 CLP & above 82,000 euros & above 4,000,001 NGN & above SG\$100,000 - SG\$999,999 US\$100,000 - US\$999,999
Table 2: Demographic answers remapped values for Annual Household Income (AHI) to handle granular responses. All available answers are listed in subsection B.1 in Appendix B.
@@ Acknowledgements
Appendix A City and image selection
Figure 6: Rankings of the selected cities for various geographic and socioeconomic attributes (i.e., average elevation, average canopy height, share of built up area in high greenness, average building height, average GDP per capita, and population density), among more than 10,000 urban centers delineated by the Global Human Settlement Urban Centre Database. (Data source: Global Human Settlement Urban Centre Database)
Figure 7: Distribution of visual complexity values for available filtered images grouped by location. Mean and selected thresholds for Abuja and the combined remaining cities are shown with vertical black lines.
Appendix B Survey design
B.1 Demographic questions
The first section of the survey consisted of 11 demographic questions. Below is the list of questions, the type of answer expected (single choice or open) and their available answers.
  1. 1.
    What is your gender? (Single choice)
    • Male
    • Female
    • Non-binary
    • Prefer not to say
  2. 2.
    What is your age group? (Single choice)
    • 21-29
    • 30-39
    • 40-49
    • Above 50
  3. 3.
    What is your nationality? (Single choice)
    • Dropdown list of countries
  4. 4.
    Which country do you currently live in? (Single choice)
    • Dropdown list of countries
  5. 5.
    Which city do you currently live in? (Single choice)
    • Dropdown list of cities based on the country chosen above
  6. 6.
    How long have you lived in the country you’re currently living in? (Single choice)
    • Less than 1 year
    • 1-3 years
    • 3-5 years
    • Above 5 years
  7. 7.
    What is your annual household income (AHI)? (Single choice)
    Singapore Netherlands USA Chile Nigeria
    \$0 - \$$9$17,999 & below \$0 - \$24,999 8,151,000 CLP & below 2,000,000 NGN & below
    \$10,000 - \$24,999 €18,000 - €29,999 \$25,000 - \$34,999 8,151,001 CLP - 20,400,000 CLP 2,000,001 NGN - 4,000,000 NGN
    \$25,000 - \$44,999 €30,000 - €47,999 \$35,000 - \$49,999 20,400,001 CLP - 40,400,000 CLP 4,000,001 NGN - 6,799,999 NGN
    \$45,000 - \$62,999 €48,000 - €64,999 \$50,000 - \$74,999 40,400,001 CLP - 60,400,000 CLP 6,800,000 NGN & above
    \$63,000 - \$81,999 €65,000 - €81,999 \$75,000 - \$99,999 60,400,001 CLP - 80,400,000 CLP Prefer not to answer
    \$82,000 - \$$9$82,000 - €$9$100,000 - \$149,999 80,400,001 CLP & above
    \$100,000 - \$119,999 €100,000 & above \$150,000 - \$199,999 Prefer not to answer
    \$120,000 - \$139,999 Prefer not to answer \$200,000 - \$999,999
    \$140,000 - \$159,999 Prefer not to answer
    \$160,000 - \$179,999
    \$180,000 - \$999,999
    Prefer not to answer
  8. 8.
    How many individuals reside in your household, including yourself? (Open numeric)
  9. 9.
    What is highest level of education? (Single choice)
    • Less than Secondary/High school
    • Secondary/High school graduate or equivalent
    • College or Vocational training or Diploma
    • Bachelor’s degree
    • Master’s degree
    • Doctorate or professional degree
    • Other
  10. 10.
    Which race or ethnicity best describes you? (Single choice)
    • Asian or Pacific Islander
    • Black or African American
    • Hispanic or Latino
    • Native American or Alaska Native
    • White or Caucasian
    • Middle Eastern or North African
    • A race/ethnicity not listed here
  11. 11.
    How well do the following statements describe your personality: I see myself as someone who …(Single choice)
    • …is reserved
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …is generally trusting
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …tends to be lazy
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …is relaxed, handles stress well
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …has few artistic interests
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …is outgoing, sociable
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …tends to find fault with others
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …does a thorough job
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …gets nervous easily
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
    • …has an active imagination
      1. (a)
        Disagree strongly
      2. (b)
        Disagree a little
      3. (c)
        Neither agree nor disagree
      4. (d)
        Agree a little
      5. (e)
        Agree strongly
B.2 Subjective indicators
The second section of the survey consisted of 10 subjective indicators: Safe, lively, wealthy, beautiful, boring, depressing, live nearby, walk, cycle, and green. To ensure participants noticed the change in questions, section breaks, i.e., blank screens stating the new set of indicator questions were added, as well as showing the indicator word in bold text in the question. The survey did not display the location the images were taken. The 50 ratings per participant consisted of five pairwise comparison questions for each of the 10 indicators, and no information regarding the location of the image was given Below is the list of questions:
  • Which place looks safer?
  • Which place looks livelier?
  • Which place looks wealthier?
  • Which place looks more beautiful?
  • Which place looks more boring?
  • Which place looks more depressing?
  • Which place looks like a place you want to live nearby?
  • Which place looks better to walk?
  • Which place looks better to cycle?
  • Which place looks greener?
Participants are able to choose the left image, right image, or an option for “Both are the same to me”.
Figure 8: Screenshot of the graphical interface a respondent saw during the pairwise comparison portion of the online survey.
Appendix C Participants’ background
Country Num. Participants Num. Nationalities Num. Cities Years living in reported city
Netherlands 200 10 1 Less than 1 year: 5 1-3 years: 5 3-5 years: 13 Above 5 years: 177
Singapore 200 8 1 Less than 1 year: 0 1-3 years: 9 3-5 years: 5 Above 5 years: 186
USA 200 17 2 Less than 1 year: 9 1-3 years: 14 3-5 years: 14 Above 5 years: 163
Chile 200 9 1 Less than 1 year: 3 1-3 years: 7 3-5 years: 8 Above 5 years: 182
Nigeria 200 1 41 Less than 1 year: 7 1-3 years: 22 3-5 years: 15 Above 5 years: 156
All
5 1,000 45 46 Less than 1 year: 24 1-3 years: 57 3-5 years: 55 Above 5 years: 864
Table 3: Summary of participants’ background, including their nationalities, the cities they reside in within each country, and the number of years they have lived in said city. As mentioned in the Method sections, while the cities where the SVI was taken were prioritized (Santiago, Amsterdam, Abuja, Singapore, and San Francisco), neighboring cities were chosen to recruit enough participants. Participants in Nigeria were recruited from the entire country, and participants in the USA were recruited from San Francisco and Santa Clara. A total of 1,000 participants, 200 from each country, were recruited.
Appendix D Q Score
The method’s score for Strength of Schedule (SOS) is referred to as the Q Score. The win (W) and loss (L) ratios of image $i$ with respect to question $u$ are defined as follows:
$W_{i,u}=\frac{w_{i,u}}{w_{i,u}+l_{i,u}+t_{i,u}},\quad L_{i,u}=\frac{l_{i,u}}{w_{i,u}+l_{i,u}+t_{i,u}}$ (1)
where $w$ is the number of times an image was selected over its paired image, $l$ is the number of times an image was not chosen over its paired image, and $t$ is the number of times an image was chosen as equal to its paired image. Using this, the score for each image $i$ and question $u$ is defined as:
$Q_{i,u}=\frac{10}{3}\left(W_{i,u}+\frac{1}{n_{i}^{w}}\sum_{j_{1}=1}^{n_{i}^{w}}W_{j_{1}u}-\frac{1}{n_{i}^{l}}\sum_{j_{2}=1}^{n_{i}^{l}}L_{j_{2}u}+1\right)$ (2)
where ${n_{i}^{w}}$ is the total number of images $i$ was preferred over, ${n_{i}^{l}}$ is the total number of images $i$ was not preferred over. The first sum extends over $j_{1}$, the images that image $i$ was preferred over, and the second sum extends over $j_{2}$, the images that were preferred over $i$.
Appendix E Extended results
Figure 9: Statistical difference of perception scores by nested demographic. Perception Q scores are calculated from ratings by participants in each nested group (combinations of ‘gender $\times$ age group $\times$ AHI’) for all (without location grouping) and each location. Locations with no significant differences in any indicator and nested demographic groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per nested group) are not shown. Significance thresholds *$p<0.05$.
Figure 10: Expanded Pearson correlation between our proposed four new indicators (live nearby, walk, cycle, and green) and the six indicators predominantly used in visual urban perception studies (safe, lively, wealthy, beautiful, boring, and depressing).
(a) live nearby
(b) walk
(a) cycle
(b) green
Figure 12: Polynomial regression analyses between the proposed perceptual indicators ((a) live nearby, (b) walk, (c)cycle, and (d) green) and the six traditional ones: safe, lively, wealthy, beautiful, boring, and depressing. Best fitting curve between linear, quadratic (normal or inverted U-shaped) and cubic, in terms of $R^{2}$, are plotted in red. Linear fitting curve is shown alongside the best fitting model with $R$ values.

### v4

@@ 4.1 Data collection
In our research, we followed a multi-criteria approach that balanced geographic representation, data availability, and cultural diversity while working within practical constraints. Our selection methodology employed a three-tier filtering process: (1) continental representation to ensure global coverage, (2) availability in established datasets (MIT PP2 [6] and NUS Global Streetscapes [52]) to enable comparative analyses, and (3) consideration of cultural, urban form, and socioeconomic diversity within these constraints. This systematic approach reduced our candidate cities considerably (Global Streetscapes includes 688 cities around the world, and PP2 only covers 56) but ensured scientific rigor and comparability with existing literature. To ensure geographic diversity (1), we included the top five largest continents by population: South America, Europe, Africa, Asia, and North America. In terms of data availability (2), in South America, only three and two cities from Brazil and Chile are present in PP2, respectively. We chose Santiago as it was the only capital among the two countries listed in both PP2 and Global Streetscapes datasets, and a better cultural representation of a Hispanic city in South America. On one hand, while the number of candidate cities in Europe, Asia, and North America is higher, we opted for cities that are commonly used in perception studies due to their unique urban form and cultural characteristics [34, 43] (3): Amsterdam, Singapore, and San Francisco. On the other hand, only a handful of cities from South Africa and Botswana in Africa are present in PP2. While we aimed to prioritize the overlap between the cities in PP2 and the Global Streetscapes datasets, we did not consider Cape Town and Johannesburg, in South Africa, as representative candidates for the continent. Unfortunately, the other candidate city, Gaborone in Botswana, presented logistical difficulties when recruiting participants. Therefore, among the top five countries based on GDP in the continent (South Africa, Egypt, Algeria, Nigeria, and Morocco) [53], we opted for Abuja in Nigeria purely based on its central location in the continent and its presence in the Global Streetscapes. This city selection, despite its limitations, provides sufficient diversity in urban morphology, cultural contexts, and socioeconomic conditions (Supplementary material Fig. 1) to test our core hypothesis about demographic influences on urban perception. We followed the image download process detailed in [52] and downloaded all available street view images within a 2.4 km2 square in the city center. For the city of Abuja, Nigeria, we manually selected a 2.4 km2 region with more urban development since the images in the Global Streetscapes dataset contained mostly highways scenes.
@@
Based on the contextual metadata available in the dataset, we selected only images with good quality, no reflection, non-panoramic, captured in clear weather, and taken during the day with front or back view direction. For image validity screening, we used visual complexity, as suggested in the Global Streetscapes dataset [52]. A low visual complexity value means the image is composed of predominantly one element, e.g., a road or wall. Supplementary material Fig. 2 shows the distribution of visual complexity values for all filtered images grouped by city. Images from all cities but Abuja have a clearer overlap in their distribution, with a visual complexity mean of 1.7; thus, we chose a conservative threshold of 1.5 and retained all images with at least this visual complexity value. However, images from Abuja have lower visual complexity values, with a mean of 1.3. While this overall lower visual complexity may reflect the intrinsic urban landscape of the selected region in the city, we aimed to have images with similar characteristics across all cities. Therefore, we retained images with at least a visual complexity of 1.3 as a trade-off for image diversity; there are twice the number of images with a visual complexity $\geq 1.3$ than with a visual complexity $\geq 1.5$.
@@ 4.2.1 Demographic survey
Participants were recruited from the general population, following their national representation by age (adults over 21 years) and gender for each country, with soft quotas for income level. The latter means that for the greater Amsterdam area, Singapore, and San Francisco and Santa Clara, participants were recruited such that their annual income ranges are close in quantity. For Chile and Nigeria, participants were recruited on a natural fallout basis. The available answers for the annual income ranges were adjusted to each country’s currency and situation based on the vendor’s suggestions, and the entire survey was translated into the official language in each country where it was deployed. Participants who did not reside in either of the five countries were excluded. A complete list of questions and their possible answers is listed in Supplementary material Section 2.
@@ 4.2.2 Human perception rating
In the online survey, each participant rated five random unique pairs of images per indicator, totaling 50 pairwise comparisons across all 10 indicators. We chose the pairwise comparison scoring method because it has been shown to produce more stable results in urban visual perception studies [34]. Supplementary material Fig. 3 shows a screenshot of the online survey for one pairwise comparison under the lively indicator. Participants clicked on the image they preferred as an answer to the indicator question, e.g., the question “Which place looks livelier” in the screenshot, or chose “Both are the same to me” for an equally rated answer. Participants were not given a predetermined definition of the perceptual indicators. Instead, we aimed to captured their broad subjective interpretation of the ten different indicators. The complete perception survey questions are listed in Supplementary material Section 2.2.
@@ 4.3 Dataset
Our final dataset consists of images showcasing urban scenes from the cities of Santiago, Amsterdam, Abuja, Singapore, and San Francisco. A total of 400 images, 80 per city, were collected, and 200 respondents from each city (or neighboring areas) were recruited as raters, adding to 1,000 respondents in total. After deploying the online survey for two months, each participant answered a total of 61 questions, consisting of 11 demographic questions and 50 pairwise comparisons. We name this dataset Street Perception Evaluation Considering Socioeconomics (SPECS) and release it openly. Table 1 shows a summary of our collected dataset in the reporting format proposed by [34] and Supplementary material Table 1 shows the demographic background of the recruited participants from each country.
@@
Raters Demographics, sample size, number of ratings per image, recruitment process We recruited 1,000 respondents (200 from each country) who reside in the city, or neighboring cities, where the images were extracted – Santiago, greater Amsterdam area, Nigeria, Singapore, and San Francisco and Santa Clara – and are above 21 years old. An equal distribution of respondents, based on annual income range, was prioritised for all but Chile and Nigeria. The gender ratio is 1:1 within countries, and the age group ratio is roughly 1:1:1:2 for 21-29, 30-39, 40-49, and above 50. More than 78% of participants from each country have lived there for more than five years, and less than 5% have lived there for less than a year. Background information from all participants is shown in Supplementary material Table 1. Each participant completed 50 ratings, with each image receiving an average of 25 ratings (min. 10, max. 46) per indicator.
Scoring Scoring methodology, rationale for method selection, evaluative indicators We selected pairwise comparison as the ranking scoring method due to its proven stability over and fairly quick assessment. On the online survey, participants were asked to choose their preference between two images given an indicator. To avoid a forced choice scenario that could lead to bias [67], and to be consistent with previous and future studies, the option for an equally rated preference (i.e., “Both are the same to me”) was shown too. We asked participants to rate a random pair of images five times for each of the 10 indicators, making up the 50 ratings per participant. Supplementary material Fig. 3 shows a screenshot of the online survey.
@@ 4.4 Data analysis
We calculated the perception scores using the Strength of Schedule (SOS) method, first used in urban perception in the Place Pulse 1.0 dataset [9]. This method’s score is referred to as the Q score, and it reflects the magnitude of the perception, bounded in $[0,10]$. Another commonly used method is the TrueSkill score [68], but unlike Q scores, it requires a higher minimum number of pairwise comparisons; i.e., 22 and 29, respectively; to reach stable scores [34]. Details about the Q score calculation are found in Supplementary material Section 4. Relative scores such as Q scores and TrueSkill scores inherently depend on the ratings considered for their calculations. Thus, rather than being an absolute measure of an image’s perceptual dimensions, these scores are only comparative; they are only valid within the rating pool in which they were calculated. As shown in the bottom panel of Figure 1, step 3, we also included the predicted perception scores for all images using an existing perception model. This perception model, a Vision Transformer (ViT) developed by [69] and used in the Global Streetscapes dataset [52], serves as a proxy of a global or one-size-fits-all approach as it was fine-tuned with the PP2 dataset [6]. Specifically, the model was trained on the top and bottom 5% rated images for each perception dimension, which were selected as the representative positive and negative samples, respectively, adapting the method by [5]. The model was then trained, with a 80%-to-20% train-test split, to predict the probability of an image being assigned to the positive label, which was subsequently multiplied by 10 to obtain the perception score. This model, which we refer to as ViT-PP2, outputs perception scores for the six traditional indicators of the PP2 dataset and reported a higher accuracy in all but the boring indicator compared to the model proposed with the PP2 dataset.
@@
Lower
20,400,000 CLP
& below
29,999 euros
& below
2,000,000 NGN
& below
SG\$0 -
SG\$44,999
US\$0 -
US\$49,999
Middle
20,400,001 CLP -
60,400,000 CLP
30,000 euros -
81,999 euros
2,000,001 NGN-
4,000,000 NGN
SG\$45,000 -
SG\$99,999
US\$50,000 -
US\$99,999
Upper
60,400,001 CLP
& above
82,000 euros
& above
4,000,001 NGN
& above
SG\$100,000 -
SG\$999,999
US\$100,000 -
US\$999,999
Table 2: Demographic answers remapped values for Annual Household Income (AHI) to handle granular responses. All available answers are listed in the Supplementary material Section 2.